OpenStudy (anonymous):

find the slop of each line y=- 2 over 5x-1 and 4x+10y=10 and determine whether they are parallel, perpendicular, or neither.

OpenStudy (anonymous):

I got y=-2/5x+1 compared to y=-2/5x-1 and that they are perpendicular to each other

OpenStudy (anonymous):

am I correct?

OpenStudy (anonymous):

If the Product of slopes of two lines =-1 , then they are perpendicular. and if slopes are equal then parallel.

OpenStudy (anonymous):

Can You write the Equation of first line again as in same as that of second line

OpenStudy (anonymous):